myininaya (myininaya):

find x and y: 19x-17y=1

OpenStudy (anonymous):

Let x or y =0

myininaya (myininaya):

no lol i'm not trying to find x and y intercepts

myininaya (myininaya):

19=17(1)+2 17=2(8)+1 => 17-2(8)=1 => 17-(19-17)8=1 => 17(9)+19(-8)=1 y=-9 x=-8

myininaya (myininaya):

oh i see what you are saying ok i forgot to say x and y cannot be zero

myininaya (myininaya):

lol so you win